Based on mass spectrometry, infrared spectroscopy, H NMR and CNMR, what could the following two unknowns be?

1.Mass Spectrometry (not shown): [M] = 134 m/z Infrared Spectroscopy (not shown): 3073, 2924, 1714, cm H Nuclear Magnetic Reso2.Mass Spectrometry (not shown): [M] = 156 (75%), [M+2] = 158 (100%), (M+4) = 160 (25%) m/z Infrared Spectroscopy (not shown):
Answer #1

1.

Applying rule of 13,

134/13 = 10 + 4/13

The base formula of the molecule = C10H(10+4) = C10H14

The peak at 1714 cm-1 in IR spectrum suggests that the molecule has a C=O moiety.

Thus replacing "CH4" unit from the base formula with O to get the molecular formula = C9H10O

Degree of unsaturation = (2c + 2 - h)/2    (where, c and h are the no. of C and H atoms in the molecule)

                                      = (2 x 9 + 2 - 10)/2

                                      = 5

The degree of unsaturation suggests that the molecule is most likely have a phenyl ring in its structure along with C=O. The assertion that the molecule has a phenyl ring can be confirmed by the multiplet peak with integration 5 at ~7.3 ppm in 1H NMR spectrum.

The two singlet with integration 3 and 2 in the NMR spectrum suggests that there are methyl and methylene moieties in the molecule and they do not have any hydrogen in the neighboring carbon(s) to couple with.

The peak at ~205 in 13C NMR indicates the presence a of ketone group in the molecule.

The peaks at 3073 cm-1 and 2924 cm-1 in IR spectrum indicates the Csp2-H and Csp3-H stretches respectively.

1H NMR peak assignment:

B СНС: О- СНС. IO в раи 5H PPM з 2H Зн

13C NMR peak assignment:

D G 220 220 200 180 160 140 120px 100 120 100 80 60 40 20 PPM

Thus the structure that satisfies the given data:

phpyhbYYl.png